OpenStudy (anonymous):

What is the solution to 3 times the square root of 4 minus 2 times the square root of 4?

OpenStudy (muzzack):

1√4

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

Ok, so the first thing you want to do is to write out what they're saying. Like this: 3 x √4 - 2√4.

OpenStudy (anonymous):

What do you think you should do next?

OpenStudy (anonymous):

If you're not sure, I'll give you one more step. Take this 3 x √4 - 2√4 and multiply 3 and √4 together. What next?

OpenStudy (anonymous):

im lost.

OpenStudy (redohawk):

the square root of four is 2

OpenStudy (redohawk):

2 * 3 = 6

OpenStudy (redohawk):

and then again the square root of 4 is 2, and 2*2 = 4

OpenStudy (anonymous):

none of these answers you guys have given me, are answers... on E2020.

OpenStudy (redohawk):

and last... 6-4 = 2

OpenStudy (redohawk):

its not 2? :(

OpenStudy (anonymous):

wait nevermind thanks redo dude

OpenStudy (anonymous):

Write 3 times the square root of 2 plus 2 times the square root of 3 in simplest form.

OpenStudy (redohawk):

oh okay, np :) gimme a medal if it is right ;) P.S. just became your fan :D

OpenStudy (redohawk):

another quesiton?